A traveler first drives 21.2 km east, then 30.6 km southeast, and finally 9.9 km south. Find the traveler's total displacement.

magnitude
direction

south of east

I got 59.354 which I put into sig figs as 59.4 km for magnitude and for direction I got 52.7367 into sig figs as 52.7 degrees which was counted incorrect
Answer #1

net horizontal component

x = 21.2 + 30.6 cos 45 = 42.837

net vertical component

y = 9.9+ 30.6 sin 45 = 31.537

net magnitude,

R^2 = x^2 + y^2

R = 53.2 km

==========

Direction

phi = arctan ( y/x)

phi = 36.36 south of east

==========

do comment in case any doubt, will reply for sure . Goodluck